# Contrasting exchange-field and spin-transfer torque driving mechanisms in all-electric electron spin resonance

### v1

@@ Appendix D Insights on mapping the QME onto a spin problem
$\langle\dot{S}_{y}\rangle$ $=$ $\left[\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg})(\rho_{ee}-\rho_{gg})+(\Delta_{ge}-2\sum_{\alpha}P_{\alpha}\cos\theta(\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})-\mbox{Im}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}))\mbox{Re}(\rho_{ge})\right.$
-- $\left.2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})\mbox{Im}(\rho_{ge})\right],$
which can be expressed independently of the singly occupied state populations and coherences by introducing $\hbar\mbox{Re}(\rho_{ge})=\langle S_{z}\rangle\sin\theta-\langle S_{x}\rangle\cos\theta$, $\langle S_{y}\rangle=-\hbar\mbox{Im}(\rho_{ge})$, and $\rho_{ee}-\rho_{gg}=2(\langle S_{z}\rangle\cos\theta+\langle S_{x}\rangle\sin\theta)/\hbar$, which follow from Eqs. 33 when $\phi=0$. This leads to:
@@
By considering that $\sum_{\alpha}P_{\alpha}\left[\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}-\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})\right]\sin\theta=\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))$, one sees that the second and third lines in the $S_{x}$ equation, both proportional to the polarization, cancel out, while in the $S_{z}$ equation, we obtain the non-equilibrium spin accumulation. At the same time, this relation further simplifies the equations for the transient state by merging the second and third terms in both equations, ultimately leading to Eqs. LABEL:spin_equations.

### v2

@@ Appendix E Insights on mapping the QME onto a spin problem
- $\left.(\Delta_{ge}-2\sum_{\alpha}P_{\alpha}(\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})-\mbox{Im}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}))\mbox{Im}(\rho_{ge})\cos\theta-2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})\mbox{Re}(\rho_{ge})\right],$
$\langle\dot{s}_{y}\rangle$ $=$ $\sum_{v=0,2}\left[\mbox{Im}(\Gamma_{ev,vg})(\rho_{ee}-\rho_{gg})-2\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})\mbox{Im}(\rho_{ge})\right]+(\Delta_{ge}-2\sum_{\alpha}P_{\alpha}(\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})-\mbox{Im}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}))\mbox{Re}(\rho_{ge})\cos\theta,$
which can be expressed independently of the singly occupied state populations and coherences by introducing $\hbar\mbox{Re}(\rho_{ge})=\langle s_{z}\rangle\sin\theta-\langle s_{x}\rangle\cos\theta$, $\langle s_{y}\rangle=-\hbar\mbox{Im}(\rho_{ge})$, and $\rho_{ee}-\rho_{gg}=2(\langle s_{z}\rangle\cos\theta+\langle s_{x}\rangle\sin\theta)/\hbar$, which follow from Eqs. (34) when $\phi=0$. This leads to:
@@
By considering that $\sum_{\alpha}P_{\alpha}\left[\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}-\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})\right]\sin\theta=\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))$, one sees that the second and third lines in the $s_{x}$ equation, both proportional to the polarization, cancel out, while in the $s_{z}$ equation, we obtain the non-equilibrium spin accumulation. At the same time, this relation further simplifies the equations for the transient state by merging the second and third terms in both equations, ultimately leading to Eqs. (17a) and (17b).